Sam is recklessly driving 100km/h in a 50km/h speed zone when he suddenly sees the police. he steps on the brakes and slows to 50km/h in three seconds. how far does he travel while braking?
1
Expert's answer
2018-10-03T06:47:08-0400
Let’s first convert km⁄h to m⁄s:
v_i=100 km/h∙(1000 m)/(1 km)∙(1 h)/(3600 s)=27.8 m/s,
v_f=50 km/h∙(1000 m)/(1 km)∙(1 h)/(3600 s)=13.9 m/s.
We can find how far does he travel while braking from the kinematic equation:
d=(v_i+v_f)/2∙t,
here, v_i is the initial velocity of Sam’s car, v_f is the final velocity of Sam’s car, t is the time during which he braking.
Then, we get:
d=(v_i+v_f)/2∙t=(27.8 m/s+13.9 m/s)/2∙3.0 s=62.5 m.